OpenStudy (anonymous):

4/5x + 5 < -3 A. x < –10 B. x ≤ 0 C. x > 5 D. x ≥ –10

OpenStudy (johnweldon1993):

Well what you want to do first....is isolate the 'x' so subtract 5 from both sides...this will leave you with \[\frac{ 4 }{ 5 }x < -8\] Now you would multiply both sides of the equation by 5/4 to solve for 'x' \[x = -8 \times \frac{ 5 }{ 4 }\] Which leaves you with \[x < -10\]